What is the time constant for the discharge of this capacitor? Enter your answer in seconds. 212> *212 Enter answer...
Answer #1

We know that Time Constant of the RC circuit

T= RC

here R=2+2 = 4 ohms ( Since 2 & 2 resistances are connected in series with capacitor ) , C=1 F

T = 4*1

T = 4seconds Answer